OpenStudy (anonymous):

Using each of the Integers (-2), (+8), (-6), and (+1) at least once, create four different expressions that include + - / and x at least once and result in an answer between (-10) and (+10) Using each of the Integers (-2), (+8), (-6), and (+1) at least once, create four different expressions that include + - / and x at least once and result in an answer between (-10) and (+10) @Mathematics

OpenStudy (eyust707):

do u just want an answer?? or do u wanna know how

OpenStudy (anonymous):

i dont know how and i need ans answer lol

OpenStudy (eyust707):

would you like to learn how?

OpenStudy (anonymous):

um well can u plz give me the equations? (answer)

OpenStudy (eyust707):

hehe you dont learn anything that way..

OpenStudy (anonymous):

ik

OpenStudy (eyust707):

sorry someone else might do your homework for you but I won't

OpenStudy (anonymous):

ok then teach me

OpenStudy (eyust707):

ok so with this one we gotta kinda anticipate what will happen.. since the numbers need to be between that rang the goal here is to keep the numbers reletively small

OpenStudy (eyust707):

also we have to make a divion so well have to use easy number for that

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

we have -2 8 -6 and 1 the division should be done first to get it out of the way

OpenStudy (anonymous):

...okk

OpenStudy (eyust707):

so divide two of those numbers

OpenStudy (anonymous):

any two?

OpenStudy (eyust707):

yep any two.. but i would pick an easy two

OpenStudy (anonymous):

like?

OpenStudy (eyust707):

thats what you gotta figure out...

OpenStudy (eyust707):

do you know what division is?

OpenStudy (anonymous):

yes of course! so i could divide 8 and -6?

OpenStudy (eyust707):

well those two would be difficult to divie because the answer is 8/6 = 4/3

OpenStudy (eyust707):

try another two

OpenStudy (anonymous):

ok, um -2 / 1?

OpenStudy (eyust707):

that works... but we need to keep our numbers small because we want our answer to be small.. so if it was me i would use this opportunity to get rid of a big number

OpenStudy (anonymous):

ok 8 / -2

OpenStudy (eyust707):

yea! thats your best bet

OpenStudy (eyust707):

so whats that equal

OpenStudy (anonymous):

-4

OpenStudy (eyust707):

ok so now we have -4.. we need to either + - or * to any of these numbers -2 -6 or 1

OpenStudy (eyust707):

you can decide how to do this next step just remember we need to keep our numbers small

OpenStudy (anonymous):

im confused already

OpenStudy (eyust707):

ok so we have -4

OpenStudy (eyust707):

and we have either three options we can add we can subtract or we can multiply

OpenStudy (eyust707):

which one would you like to do

OpenStudy (anonymous):

for what?

OpenStudy (eyust707):

for the next step

OpenStudy (anonymous):

....so what options do i have? i dont know what to do

OpenStudy (eyust707):

ok it seems you dont understnad what the question is asking for

OpenStudy (eyust707):

lets look back up to the top

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

Using each of the Integers (-2), (+8), (-6), and (+1) at least once, create four different expressions that include + - / and x at least once and result in an answer between (-10) and (+10) this seems confusing but i promise its hecka easy once you it clicks

OpenStudy (anonymous):

ok whatever u say!

OpenStudy (eyust707):

ok so basically what they want you to do is add subtract multiply and divide

OpenStudy (anonymous):

right

OpenStudy (eyust707):

and you must you those number they give you once each

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

the only restrictoin is that your number is smaller than 10

OpenStudy (anonymous):

yup

OpenStudy (eyust707):

otherwise you are free to add sub mult and divide those numbers in any sequence you want

OpenStudy (anonymous):

m hm

OpenStudy (eyust707):

we divided first cause it usually is the trickiest

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

but to satisfy the question we still need to multiply add and subtract

OpenStudy (anonymous):

ok so can i multiply next then?

OpenStudy (eyust707):

sure! =)

OpenStudy (anonymous):

ok, but what do i multiply?

OpenStudy (eyust707):

so we ended with -4 what would you like to multiply it by?

OpenStudy (anonymous):

so out of the given numbers right?

OpenStudy (eyust707):

yeo but we cant use one we already used

OpenStudy (anonymous):

ok so multiply -4 by . . . -6?

OpenStudy (eyust707):

well we want small numbers... rember our restricton

OpenStudy (anonymous):

ok so multiply by .... what would be the best?

OpenStudy (eyust707):

lol theres only one other option

OpenStudy (anonymous):

....whic is? . . . -2?

OpenStudy (anonymous):

no wait +1

OpenStudy (eyust707):

we used negitive 2 in the first exquation

OpenStudy (eyust707):

yep!

OpenStudy (anonymous):

ok so then go -4 x +1

OpenStudy (eyust707):

ok so so far we have 8/-2 = -4 -4 * 1 =

OpenStudy (anonymous):

-4

OpenStudy (eyust707):

yeo whats it equal

OpenStudy (eyust707):

=)

OpenStudy (eyust707):

ok do you want to plus or minus?

OpenStudy (anonymous):

umm plus .. .-6?

OpenStudy (eyust707):

sure

OpenStudy (anonymous):

which equals -5, and the i have to subtract . . but i used all the numbers

OpenStudy (eyust707):

8/-2 = -4 -4 * 1 = -4 -4 + -6 = ??

OpenStudy (anonymous):

-5

OpenStudy (eyust707):

-4 + -6 is the same as -4 - 6

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

what -1 -1 = ??

OpenStudy (anonymous):

-2

OpenStudy (eyust707):

ok.. what about -2 -2 =??

OpenStudy (anonymous):

-4

OpenStudy (eyust707):

nice!

OpenStudy (eyust707):

what about -4-6=??

OpenStudy (anonymous):

-10 .. why r u asking me these? arent we doing the equation

OpenStudy (eyust707):

nice! hehe because you just told me that -4 -6 = -5 ; )

OpenStudy (anonymous):

ooo

OpenStudy (eyust707):

8/-2 = -4 -4 * 1 = -4 -4 + -6 = -10 -10 -?? = ???

OpenStudy (anonymous):

7

OpenStudy (eyust707):

ok so if you read it you have to use each number "at least once" this means you are allowed to use a any number twice if you would like

OpenStudy (anonymous):

ok...

OpenStudy (eyust707):

so pick one that will make -10 become closer to 0

OpenStudy (anonymous):

umm.... -6?

OpenStudy (eyust707):

ding ding ding!!

OpenStudy (eyust707):

8/-2 = -4 -4 * 1 = -4 -4 + -6 = -10 -10 -(-6) = ???

OpenStudy (anonymous):

yay! lol so i subtract that from 7?